OpenStudy (anonymous):

The height above ground (feet) of a ball thrown straight upward is given by h(t)=3.5+8t-16t^2, where t is in seconds. a) What is the average velocity of the ball during the first half second? b) What is the ball's velocity at t=1/2? c) What is the ball's acceleration at t=1/2?

OpenStudy (owlcoffee):

Let's first take when it initially is in "rest", with that I mean t=0 So replacing: \[h:h(t)=3.5+8t-16t ^{2} \] si the image when t=0 is: \[h:h(0)= 3.5+8(0)-16(0)^{2}=3.5\]

OpenStudy (anonymous):

ok

OpenStudy (dan815):

average velocity is the slope of the line draw from time =0 to time = 1/2

OpenStudy (dan815):

for example your function looks something like this

OpenStudy (owlcoffee):

Now you want to know in what's the image of h when t=1/2, can you do that?

OpenStudy (dan815):